User Tools

Site Tools


lbaops:lbamay2010:v253apalog

Recording to ATNF V014B.

Pols incorrect. Wrong feed installed so we have linear rather than circulars. After discussions between Chris and the PI, it was decided that changing the feed was the best course of action. Recording stopped at 11:12:00 UT, dish stowed prior to that to allow feed change.

Valid data from just prior to 12:27:00UT. Pols are now circular but are SWAPPED. Will swap cables after this experiment.

Sources set at Parkes at arround 19:22:30UT. Stopped recorder at about 19:26UT.

lbaops/lbamay2010/v253apalog.txt · Last modified: 2015/12/18 16:38 (external edit)